# Spectral Moment Formulae for $\hbox{GL}(3)\times\hbox{GL}(2)$ L-functions II: The Eisenstein Case

### v2

@@ 2. Notations and conventions
We fix $\eta>40$. Our test function $H$ lies in the class $\mathcal{C}_{\eta}$. The function $h$ is the Kontorovich–Lebedev inversion of $H$; see (3.13).
@@ 2.5. Measures
$d^{W}\mu\,:=\,|\Gamma(\mu)|^{-2}\,\frac{d\mu}{4\pi i}\hskip 20.0pt(\mu\,\in\,i\mathbb{R}).$ (2.2)
@@
2.6. Analysis
The Euler $\Gamma$-function, i.e., the analytic continuation of $\int_{0}^{\infty}e^{-x}x^{s}\,d^{\times}x$, is denoted by $\Gamma(s)$; and $\Gamma_{\mathbb{R}}(s):=\pi^{-s/2}\Gamma(s/2)$, for $s\in\mathbb{C}$. We use both $\Gamma(s)$ and $\Gamma_{\mathbb{R}}(s)$ as the former is common in our references, but the latter simplifies our formulae on several occasions.
The contours of the Barnes integrals $\int_{-i\infty}^{i\infty}\,\cdots\,\int_{-i\infty}^{i\infty}\ (\cdots)\,\frac{ds_{1}}{2\pi i}\,\cdots\,\frac{ds_{k}}{2\pi i}$ are chosen so that they pass to the right of all poles of the gamma functions of the form $\Gamma(s_{i}+a)$, and to the left of all poles of those of the form $\Gamma(a-s_{i})$.
Let $\Lambda(s):=\Gamma_{\mathbb{R}}(s)\zeta(s)$ denote the completed Riemann zeta function. It admits a holomorphic continuation to $\mathbb{C}$ except for simple poles at $s=0,1$, and satisfies the functional equation $\Lambda(s)=\Lambda(1-s)$ for any $s\in\mathbb{C}$.
The $K$-Bessel function is defined by
for any $y>0$ and $\mu\in\mathbb{C}$.
The Gauss $_{2}F_{1}$ hypergeometric function is defined by
provided that both integrals converge absolutely.
@@ 2.7. Automorphic objects
The Fourier–Whittaker coefficients and periods of an automorphic form $\phi$ of $\Gamma_{2}$ are denoted, respectively, by $\mathcal{B}_{\phi}(a)$ and $\mathcal{W}_{a}(g;\,\phi)$; those of an automorphic form $\Phi$ of $\Gamma_{3}$ are denoted by $\mathcal{B}_{\Phi}(m_{1},m_{2})$ and $\mathcal{W}_{(m_{1},m_{2})}(g;\,\Phi)$. We write $W_{\mu}(g)$ and $W_{\alpha}(g)$ for the standard Whittaker functions for $\overline{\mathrm{G}}_{2}$ and $\overline{\mathrm{G}}_{3}$ respectively, and $W_{\alpha,\,w}^{(0,1)}(\mathbf{y})$ and $W_{\alpha,\,w}^{(1,0)}(\mathbf{y})$ for the degenerate Whittaker functions for $\overline{\mathrm{G}}_{3}$. This is no ambiguity as the group to which $g$ belongs is clear from the context. We use $\mu$ and $\alpha$ to denote the spectral parameters for, respectively, the automorphic forms of $\Gamma_{2}$ and $\Gamma_{3}$ (or Whittaker functions). In our case, $\mu$ is always purely imaginary because of (2.2) and the known Selberg’s conjecture for $\Gamma_{2}$ ([Gold, Theorem 3.7.2]).
@@
whenever the integral converges absolutely. In our previous works [Kw23, Kw25], the integral (2.8) was instead denoted by $(\phi,\ (\mathbb{P}_{2}^{3}\widetilde{F})\,|\det*|^{\overline{s}-\frac{1}{2}})_{\Gamma_{2}\setminus\mathrm{GL}_{2}(\mathbb{R})}$, where $\widetilde{F}(g):=F(^{t}g^{-1})$ for $g\in\mathrm{G}_{3}$.
@@ 3. Preliminary
In this section, we collect results from the sources [Gold, Bump84] that are essential to our arguments. We have adjusted the conventions to strive for consistency and to follow the recent shift of conventions (closest to [Bu20, Bu18]), which better align with those in the theory of automorphic representation. Part of the preliminary results can also be found in [Kw23].
@@ 3.1. Barnes’ identities
See [Ba64, Chapter 6.2]. This is known as the second Barnes lemma. ∎
@@ 3.2. Whittaker functions and transforms
The integral of (3.6) converges absolutely whenever $\operatorname{Re}(\alpha_{1}-\alpha_{2})>0$ and $\operatorname{Re}(\alpha_{2}-\alpha_{3})>0$, and $W_{\alpha}^{\pm}$ admits a holomorphic continuation to $\mathfrak{a}_{\mathbb{C}}$; see [Gold, Chapter 5.5]. Several normalizations of the standard Whittaker function are present in the literature. We opt for the one such that the functional equation $W_{\alpha}^{\pm}(g)=W_{\alpha^{w}}^{\pm}(g)$ holds for any $w\in W_{3}$ and $g\in\mathrm{G}_{3}$.
@@
The sign convention of the $\alpha_{i}$’s in (3.7) is consistent with [Bu20], but is opposite to that of [Gold, eqs. (6.1.4) & (6.1.5)]. It is important to keep track of this in Lemma 3.7, Propositions 6.146.15, and Section 6.8, where the symmetry must be matched judiciously with the CFKRS conjectures.
@@
Follows directly from Lemma 3.4, contour shifting, and Stirling’s formula. Sharper and more uniform bounds can be found in [Blo13, Proposition 1] and [BHM20, Theorem 1]. ∎
We will need the explicit evaluation of the archimedean $\hbox{GL}(3)\times\hbox{GL}(2)$ Rankin–Selberg integral, which is a consequence of the second Barnes Lemma.
@@
This was proved in [Bump88]. To clarify our conventions, we provide a short proof here. Take $\sigma_{1}\in(\operatorname{\bm{\alpha}}+|\operatorname{Re}\mu|,\,\sigma)$. From Lemma 3.4 and Mellin inversion, we have, for $\operatorname{Re}s_{0}>\operatorname{\bm{\alpha}}$, that
@@
From [Mo97, Lemma 2.6], one also has $(h^{\#})^{\flat}(y)=h(y)$ for $h\in C_{c}^{\infty}(0,\infty)$.
@@ 3.3.1. $\mathrm{PGL}(2)$
The invariant differential operator on $\mathfrak{h}^{2}$ is the hyperbolic Laplacian $\Delta:=-y^{2}(\partial_{x}^{2}+\partial_{y}^{2})$. An automorphic form $\phi:\mathfrak{h}^{2}\rightarrow\mathbb{C}$ of $\Gamma_{2}=\mathrm{SL}_{2}(\mathbb{Z})$ satisfies $\Delta\phi=(1/4-\mu^{2})\phi$ for some $\mu=\mu(\phi)\in\mathbb{C}$. The Fourier coefficient of $\phi$, denoted by $\mathcal{B}_{\phi}(a)$, is defined, for any $a\in\mathbb{Z}-\{0\}$, by
$\mathcal{W}_{a}(y;\phi)\ :=\ \int_{\mathbb{Z}\setminus\mathbb{R}}\ \phi\left[\begin{pmatrix}1&x\\ &1\end{pmatrix}\begin{pmatrix}y&\\ &1\end{pmatrix}\right]e(-ax)\,dx\ =\ \frac{\mathcal{B}_{\phi}(a)}{\sqrt{|a|}}W_{\mu(\phi)}(|a|y).$ (3.17)
Let $I_{\mu}(y):=y^{\mu+\frac{1}{2}}$, which is also a function on $\mathrm{U}_{2}\setminus\mathfrak{h}^{2}$. The Eisenstein series of $\Gamma_{2}$ is defined by
@@
The series (3.18) converges absolutely for $\operatorname{Re}\mu>1/2$ and admits a meromorphic continuation to $\mathbb{C}$; see [Gold, Chapter 3.1]. We have $\Delta E(*;\mu)=(1/4-\mu^{2})E(*;\mu)$, and the Fourier coefficients of $E(*;\mu)$ are given by
@@ 3.3.2. $\mathrm{PGL}(3)$
Let $m=(m_{1},m_{2})\in(\mathbb{Z}-\{0\})^{2}$ and $\Phi:\Gamma_{3}\setminus\mathfrak{h}^{3}\rightarrow\mathbb{C}$ be an automorphic form. The $(m_{1},m_{2})$-th Fourier–Whittaker period of $\Phi$ refers to the integral
@@
The $(m_{1},m_{2})$-th Fourier coefficient of $\Phi$ refers to the complex number $\mathcal{B}_{\Phi}(m_{1},m_{2})$ such that
@@
As explained in [Kw23, Section 4], it is equivalent to consider the unipotent period
@@
for $\Psi:=\rho(w_{\ell})\Phi$. Recall the following construction used in the integral representation of the $\mathrm{GL}_{3}$ standard $L$-function (see [Gold, Chapters 6.5 & 12.3]):
@@
Due to (3.20) and $(\mathbb{P}_{1}^{3}\Psi)(g)\,=\,\sum_{n\neq 0}\,\mathcal{W}_{(1,n)}(g;\,\Psi)$, observe that upon inserting the second term on the right-hand side of (3.27) into our Fourier–Hecke period ((1.21) or (6.7)) and applying Mellin inversion, one obtains an integral moment of a product of the $\zeta$-function (from the $a$-sum) and the $L$-function of $\Phi$ (from the $n$-sum; see (3.39)), with the integral transform given in (4.1). These steps were performed carefully in our previous works [Kw23, Kw25] and will be reviewed in Sections 6.5 and 6.7.
@@ 3.3.3. Eisenstein series for $\mathrm{PGL}(3)$
We require explicit computations of the Fourier–Whittaker periods for the Eisenstein series of $\Gamma_{3}$ in order to package the moment $\mathfrak{M}_{\alpha}^{(3)}(s;H)$ as a period, extract the main terms and recast them according to the CFKRS conjectures, and carry out the analytic continuation argument of Section 6.
A standard reference is [Bump84, Chapter VII]; see especially Theorem 7.2. For normalizations of the degenerate Whittaker functions of $\overline{\mathrm{G}}_{3}$, readers should refer to [Bump84, Chapter III]. It is also helpful to consult the reformulation by [Bu18, Section 4], which is more streamlined and up-to-date. Readers should beware that [Bu18] uses the “incomplete” Whittaker functions, whereas both this article and [Bump84] use the “complete” ones.
@@
The series (3.28) converges absolutely on $\big\{\alpha\in\mathfrak{a}_{\mathbb{C}}:\,\operatorname{Re}(\alpha_{1}-\alpha_{2})>1,\,\operatorname{Re}(\alpha_{2}-\alpha_{3})>1\big\}$ (see [Bu18, Section 4.1]). Its meromorphic continuation to $\mathfrak{a}_{\mathbb{C}}$ and functional equation are established explicitly in [Bump84, Chapter 8] and [Bu18]. Langlands proved these results in great generality.
We must keep track of the degenerate terms in the Fourier expansion of $E_{\min}^{(3)}(g;\alpha)$ for our applications. We prefer to apply the more compact form of the expansion (Lemma 3.13), rather than the fully explicit form used in [Li14], because it better detects:
@@
Recall the Weyl actions on the spectral parameters (see (1.4)). Translating [Bump84, eqs. (3.10)–(3.15), (3.40) & (3.45)] into the notations of this article, we have
@@
where $w\in W_{3}$, $\alpha\in\mathfrak{a}_{\mathbb{C}}$, $\mathbf{y}\in\mathrm{Y}^{+}$, and $W_{(\alpha_{1}^{w}-\alpha_{2}^{w})/2}$, $W_{(\alpha_{2}^{w}-\alpha_{3}^{w})/2}$ are standard Whittaker functions of $\overline{\mathrm{G}}_{2}$.
@@
The completion factors in (3.31) simplify many formulae and ensure that the Hecke combinatorics of the non-degenerate Fourier coefficients coincide in both the cuspidal and Eisenstein cases. They also obviate the need to consider the zero-free region of the $\zeta$-function in our argument. (Note: the formulae stated in [Li14] and [Bu18] concern the incomplete Eisenstein series.)

### v3

@@ 2. Notations and conventions
We fix $\eta>40$. Our test function $H$ lies in the class $\mathcal{C}_{\eta}$. The function $h=H^{\flat}$ is the Kontorovich–Lebedev inversion of $H$ as defined in (3.13).
2.3. Analysis
The Euler $\Gamma$-function is defined by
$\Gamma(s)\,:=\,\int_{0}^{\infty}e^{-x}x^{s}\,d^{\times}x$ (2.1)
on $\operatorname{Re}s>0$. It admits a meromorphic continuation to $\mathbb{C}$. Let $\Gamma_{\mathbb{R}}(s):=\pi^{-s/2}\Gamma(s/2)$ for $s\in\mathbb{C}$. We use both $\Gamma(s)$ and $\Gamma_{\mathbb{R}}(s)$ as the former is common in our references, but the latter simplifies our formulae on several occasions.
The contours of the Barnes integrals
$\int_{-i\infty}^{i\infty}\,\cdots\,\int_{-i\infty}^{i\infty}\ (\cdots)\ \frac{ds_{1}}{2\pi i}\,\cdots\,\frac{ds_{k}}{2\pi i}$ (2.2)
are chosen so that they pass to the right of all poles of the gamma functions of the form $\Gamma(s_{i}+a)$, and to the left of all poles of those of the form $\Gamma(a-s_{i})$.
Let $\Lambda(s):=\Gamma_{\mathbb{R}}(s)\zeta(s)$ denote the completed Riemann zeta function. It admits a holomorphic continuation to $\mathbb{C}$ except for simple poles at $s=0,1$, and satisfies the functional equation
$\Lambda(s)\ =\ \Lambda(1-s)$ (2.3)
for any $s\in\mathbb{C}$.
The $K$-Bessel function is defined by
for any $y>0$ and $\mu\in\mathbb{C}$.
The Gauss $_{2}F_{1}$ hypergeometric function is defined by
provided that both integrals converge absolutely.
@@ 2.6. Measures
$d^{W}\mu\,:=\,|\Gamma_{\mathbb{R}}(\mu)|^{-2}\,\frac{d\mu}{4\pi i}\hskip 20.0pt(\mu\,\in\,i\mathbb{R}).$ (2.8)
@@ 2.7. Automorphic objects
The Fourier–Whittaker coefficients and periods of an automorphic form $\phi$ of $\Gamma_{2}$ are denoted, respectively, by $\mathcal{B}_{\phi}(a)$ and $\mathcal{W}_{a}(g;\,\phi)$; those of an automorphic form $\Phi$ of $\Gamma_{3}$ are denoted by $\mathcal{B}_{\Phi}(m_{1},m_{2})$ and $\mathcal{W}_{(m_{1},m_{2})}(g;\,\Phi)$. We write $W_{\mu}(g)$ and $W_{\alpha}(g)$ for the standard Whittaker functions for $\overline{\mathrm{G}}_{2}$ and $\overline{\mathrm{G}}_{3}$ respectively, and $W_{\alpha,\,w}^{(0,1)}(\mathbf{y})$ and $W_{\alpha,\,w}^{(1,0)}(\mathbf{y})$ for the degenerate Whittaker functions for $\overline{\mathrm{G}}_{3}$. This is no ambiguity as the group to which $g$ belongs is clear from the context. We use $\mu$ and $\alpha$ to denote the spectral parameters for, respectively, the automorphic forms of $\Gamma_{2}$ and $\Gamma_{3}$ (or Whittaker functions). In our case, $\mu$ is always purely imaginary because of (2.8) and the known Selberg’s conjecture for $\Gamma_{2}$ ([31, Theorem 3.7.2]).
@@
whenever the integral converges absolutely. In our previous works [49, 50], the integral (2.11) was instead denoted by $(\phi,\ (\mathbb{P}_{2}^{3}\widetilde{F})\,|\det*|^{\overline{s}-\frac{1}{2}})_{\Gamma_{2}\setminus\mathrm{GL}_{2}(\mathbb{R})}$, where $\widetilde{F}(g):=F(^{t}g^{-1})$ for $g\in\mathrm{G}_{3}$.
@@ 3. Preliminary
In this section, we collect results from the sources [31, 13] that are essential to our arguments. We have adjusted the conventions to strive for consistency and to follow the recent shift of conventions (closest to [16, 15]), which better align with those in the theory of automorphic representation. Part of the preliminary results can also be found in [49].
@@ 3.1. Barnes’ identities
See [1, Chapter 6.2]. This is known as the second Barnes lemma. ∎
@@ 3.2. Whittaker functions and transforms
The integral of (3.6) converges absolutely whenever $\operatorname{Re}(\alpha_{1}-\alpha_{2})>0$ and $\operatorname{Re}(\alpha_{2}-\alpha_{3})>0$, and $W_{\alpha}^{\pm}$ admits a holomorphic continuation to $\mathfrak{a}_{\mathbb{C}}$; see [31, Chapter 5.5]. Several normalizations of the standard Whittaker function are present in the literature. We opt for the one such that the functional equation $W_{\alpha}^{\pm}(g)=W_{\alpha^{w}}^{\pm}(g)$ holds for any $w\in W_{3}$ and $g\in\mathrm{G}_{3}$.
@@
The sign convention of the $\alpha_{i}$’s in (3.7) is consistent with [16], but is opposite to that of [31, eqs. (6.1.4) & (6.1.5)]. It is important to keep track of this in Lemma 3.7, Propositions 6.146.15, and Section 6.8, where the symmetry must be matched judiciously with the CFKRS conjectures.
@@
Follows directly from Lemma 3.4, contour shifting, and Stirling’s formula. Sharper and more uniform bounds can be found in [11, Proposition 1] and [6, Theorem 1]. ∎
We will need the explicit evaluation of the archimedean $\hbox{GL}(3)\times\hbox{GL}(2)$ Rankin–Selberg integral, which is a consequence of the second Barnes Lemma. This calculation also verifies the sign convention of (3.7).
@@
This was proved in [14]. To clarify our conventions, we provide a short proof here. Take $\sigma_{1}\in(\operatorname{\bm{\alpha}}+|\operatorname{Re}\mu|,\,\sigma)$. From Lemma 3.4 and Mellin inversion, we have, for $\operatorname{Re}s_{0}>\operatorname{\bm{\alpha}}$, that
@@
From [60, Lemma 2.6], one also has $(h^{\#})^{\flat}(y)=h(y)$ for $h\in C_{c}^{\infty}(0,\infty)$.
@@ 3.3.1. $\mathrm{PGL}(2)$
The invariant differential operator on $\mathfrak{h}^{2}$ is the hyperbolic Laplacian $\Delta:=-y^{2}(\partial_{x}^{2}+\partial_{y}^{2})$. An automorphic form $\phi:\mathfrak{h}^{2}\rightarrow\mathbb{C}$ of $\Gamma_{2}=\mathrm{SL}_{2}(\mathbb{Z})$ satisfies $\Delta\phi=(1/4-\mu^{2})\phi$ for some $\mu=\mu(\phi)\in\mathbb{C}$. The Fourier coefficient of $\phi$, denoted by $\mathcal{B}_{\phi}(a)$, is defined by
$\mathcal{W}_{a}(y;\phi)\ :=\ \int_{\mathbb{Z}\setminus\mathbb{R}}\ \phi\left[\begin{pmatrix}1&x\\ &1\end{pmatrix}\begin{pmatrix}y&\\ &1\end{pmatrix}\right]e(-ax)\,dx\ =\ \frac{\mathcal{B}_{\phi}(a)}{\sqrt{|a|}}W_{\mu(\phi)}(|a|y)$ (3.17)
for any $a\in\mathbb{Z}-\{0\}$ and $y>0$.
Let $I_{\mu}(y):=y^{\mu+\frac{1}{2}}$, which can be regarded as a function on $\mathrm{U}_{2}\setminus\mathfrak{h}^{2}$. The Eisenstein series of $\Gamma_{2}$ is defined by
@@
The series (3.18) converges absolutely for $\operatorname{Re}\mu>1/2$ and admits a meromorphic continuation to $\mathbb{C}$; see [31, Chapter 3.1]. We have $\Delta E(*;\mu)=(1/4-\mu^{2})E(*;\mu)$, and the Fourier coefficients of $E(*;\mu)$ are explicitly given by
@@ 3.3.2. $\mathrm{PGL}(3)$
Let $(m_{1},m_{2})\in\mathbb{Z}^{2}$ and $\Phi:\Gamma_{3}\setminus\mathfrak{h}^{3}\rightarrow\mathbb{C}$ be an automorphic form. The $(m_{1},m_{2})$-th Fourier–Whittaker period of $\Phi$ refers to the integral
@@
When $(m_{1},m_{2})\in(\mathbb{Z}-\{0\})^{2}$, the $(m_{1},m_{2})$-th Fourier coefficient of $\Phi$ refers to the complex number $\mathcal{B}_{\Phi}(m_{1},m_{2})$ such that
@@
As explained in [49, Section 4], it is equivalent to consider the unipotent period
@@
for $\Psi:=\rho(w_{\ell})\Phi$. Recall the following construction used in the integral representation of the $\mathrm{GL}_{3}$ standard $L$-function (see [31, Chapters 6.5 & 12.3]):
@@
Due to (3.20) and $(\mathbb{P}_{1}^{3}\Psi)(g)\,=\,\sum_{n\neq 0}\,\mathcal{W}_{(1,n)}(g;\,\Psi)$, observe that upon inserting the second term on the right-hand side of (3.27) into our Fourier–Hecke period ((1.18) or (6.7)) and applying Mellin inversion, one obtains an integral moment of a product of the $\zeta$-function (from the $a$-sum) and the $L$-function of $\Phi$ (from the $n$-sum; see (3.38)), with the integral transform given in (4.1). These steps were performed carefully in our previous works [49, 50] and will be reviewed in Sections 6.5 and 6.7.
@@ 3.3.3. Eisenstein series for $\mathrm{PGL}(3)$
We require explicit computations of the Fourier–Whittaker periods for the Eisenstein series of $\Gamma_{3}$ in order to package the moment $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$ as a period, extract the main terms and recast them according to the CFKRS conjectures, and carry out the analytic continuation argument of Section 6.
A standard reference is [13, Chapter VII]; see especially Theorem 7.2. For normalizations of the degenerate Whittaker functions of $\overline{\mathrm{G}}_{3}$, readers should refer to [13, Chapter III]. It is also helpful to consult the reformulation by [15, Section 4], which is more streamlined and up-to-date. Readers should beware that [15] uses the “incomplete” Whittaker functions, whereas both this article and [13] use the “complete” ones.
@@
The series (3.28) converges absolutely on $\big\{\alpha\in\mathfrak{a}_{\mathbb{C}}:\,\operatorname{Re}(\alpha_{1}-\alpha_{2})>1,\,\operatorname{Re}(\alpha_{2}-\alpha_{3})>1\big\}$ (see [15, Section 4.1]). Its meromorphic continuation to $\mathfrak{a}_{\mathbb{C}}$ and functional equation are established explicitly in [13, Chapter 8] and [15]. Langlands proved these results in great generality.
We must keep track of the degenerate terms in the Fourier expansion of $E_{\min}^{(3)}(g;\alpha)$ for our applications. We prefer to apply the more compact form of the expansion (Lemma 3.13), rather than the fully explicit form used in [53], because it better detects:
@@
Recall that the Weyl group $W_{3}$ of $\hbox{GL}(3)$ consists of the permutation matrices
The actions of $w\in W_{3}$ on $\mathfrak{a}_{\mathbb{C}}$ are given by:
Translating [13, eqs. (3.10)–(3.15), (3.40) & (3.45)] into the notations of this article, we have
@@
where $w\in W_{3}$, $\alpha\in\mathfrak{a}_{\mathbb{C}}$, $\mathbf{y}\in\mathrm{Y}^{+}$, and $W_{(\alpha_{i}^{w}-\alpha_{i+1}^{w})/2}$ ($i=1,2$) are the standard Whittaker functions of $\overline{\mathrm{G}}_{2}$.
@@
The completion factors in (3.33) simplify many formulae and ensure that the Hecke combinatorics of the non-degenerate Fourier coefficients coincide in both the cuspidal and Eisenstein cases. They also obviate the need to consider the zero-free region of the $\zeta$-function in our argument. (Note: the formulae stated in [53] and [15] concern the incomplete Eisenstein series.)